# Understanding Oversmoothing in Diffusion-Based GNNs From the Perspective of Operator Semigroup Theory

### v1

@@ 4.1 Operator Representation of Feature
$h_{i}(u,t)=P_{t}f_{i}(u).$ (8)
The findings stand in one dimension can naturally extend to multiple dimensions, leading to:
$\begin{cases}h(u,t)=P_{t}f(u),\\ f(u)=X(u),\end{cases}\quad\text{for }u\in\mathcal{V},$ (9)
The advantage of this form becomes evident in Section 4.2, as it offers an explanation for the oversmoothing problem through ergodicity.
4.2 Ergodicity and Oversmoothing
@@
$\sum_{u\in\mathcal{V}}\mathcal{L}f(u)\pi(u)=0.$ (10)
We then define the ergodicity of operators as follows:
Definition 4.2 (Ergodicity of operator $\mathcal{L}$).
The operator $\mathcal{L}$ is said to be ergodic if every $f$ such that $\mathcal{L}f=0$ is constant.
Next, we proceed to elaborate in the context of graph diffusion. It can be verified that the generator of graph diffusion $\mathcal{A}$ exhibits ergodicity, leading to the following theorem.
Theorem 4.3.
@@
See proofs in Appendix A.2. From this theorem, one can deduce that the ergodicity of operator is dependent on the topology property of the graph. Notice that $10\leq a(u,v)<$ 1, from Perron-Frobenius theorem (Theorem 8.4.4 of Horn & Johnson (2012)), the invariant measure $\mu$ of $\mathcal{A}$ exists and is unique (Lemma A.2). Then the ergodicity of operator $\mathcal{A}$ ensures $P_{t}f(u)$ converges to the same constant for every $u\in\mathcal{V}$, as time $t\rightarrow\infty$. This finding is precisely the cause of oversmoothing as different nodes share the same convergence. Under the same assumption in Theorem 4.3, the aforementioned discussion can be summarized into the following theorem.
Theorem 4.4.
@@
See proofs in Appendix A.3. Substituting the notation from Eq. (9) in conjunction with Theorem 4.4, we can derive the following convergence:
$h(u,t)\rightarrow\sum_{v\in\mathcal{V}}X_{v}\mu(v)$ (12)
This furnishes a fundamental mathematical explanation for the occurrence of oversmoothing in diffusion-based GNNs. The underlying reasons can be attributed to the ergodicity of the operator. Consequently, to mitigate oversmoothing, it is natural to contemplate the disruption of ergodicity, as shown in the subsequent section.
4.3 Ergodicity Breaking Term
We aim to mitigate the issue of oversmoothing by disrupting the ergodicity of the operator. This necessitates a modification of the operator $\mathcal{A}$ as presented in Eq. (6). The central concept underpinning our approach involves introducing an additional term to the operator $\mathcal{A}$, yielding a new operator:
$\tilde{\mathcal{A}}:=\mathcal{A}+\mathcal{C},$ (13)
@@ 4.3 Ergodicity Breaking Term
See proofs in Appendix A.4. By substituting the operator $\mathcal{A}$ with the operator $\tilde{\mathcal{A}}$, we obtain a variation of the diffusion process (3):
$\frac{\partial h(u,t)}{\partial t}:=\mathcal{A}h(u,t)+\mathcal{C}h(u,t),$
serving as a framework for designing GNNs to mitigate oversmoothing.
4.4 Connection to Existing Work
Graph diffusion with ergodicity-breaking term $\mathcal{C}$ is an explanatory framework for many existing diffusion-based GNNs that have been proven effective in addressing the oversmoothing problem. In previous attempts to mitigate the oversmoothing issue in diffusion-based GNNs, researchers have proposed several approaches. Chamberlain et al. (2021b) adds a source term in the graph diffusion equation, while Wang et al. (2023) alleviates the issue of oversmoothing by introducing an Allen-Cahn term. Based on the reaction-diffusion equations, Choi et al. (2023) introduces the reaction term $r(H(t))$, encompassing the forms of the previous two approaches.
In comparison to the aforementioned methods (Thorpe et al., 2022; Choi et al., 2023; Wang et al., 2023), our framework offers a more rigorous theoretical foundation and provides the genetic condition to address oversmoothing. Then the three methods mentioned earlier, which address the oversmoothing issue by incorporating additional terms, can be viewed as specific cases of our approach.
5 Probability View
@@
5.1 Markov Chain on Graph with Generator $\mathcal{A}$
Given a Markov process denoted as $\xi$ characterized by its state space $E$ and a transition kernel denoted as $p(t,x,\mathrm{d}y)$, we can naturally establish a Markov semigroup as follows:
$P_{t}f(x):=\int_{E}f(y)p(t,x,\mathrm{d}y)=\mathbb{E}^{x}[f(\xi_{t})],$
where $\mathbb{E}^{x}[\cdot]=\mathbb{E}[\cdot|\xi_{0}=x]$ represents the expectation operator under the probability measure $p(\cdot,x,\mathrm{d}y)$. The generator of the Markov process $\xi$ can be obtained via Equation (5). Conversely, in accordance with the Kolmogorov theorem (Theorem 2.11 of (Ge & Blumenthal, 2011)), there exists a Markov process given its transition function and initial distribution. In this subsection, commencing from the generator $\mathcal{A}$ of graph diffusion, we formulate the corresponding process denoted as $\mathbf{X}$, which serves as a representation for the node features $H(t)$ on the graph.
Let $(\Omega,\mathscr{F},\mathbb{P})$ be a probability space and $\mathbf{Y}=\{\mathbf{Y}_{n}\}_{n\in\mathbb{N}}$ be a Markov chain with state space $(\mathcal{V},\mathscr{S})$ such that
$\mathbb{P}(\mathbf{Y}_{n+1}=v|\mathbf{Y}_{n}=u)=a(u,v).$
The jumping times are defined as:
@@ 5.1 Markov Chain on Graph with Generator $\mathcal{A}$
where $\zeta_{k},\;k=1,2,\ldots,n$ is a sequence of independent exponentially distributed random variables of parameter $1$. Next, we introduce the right-continuous process $\mathbf{X}:[0,\infty)\times\Omega\rightarrow X$ defined as:
$\mathbf{X}_{t}:=\mathbf{Y}_{n},\;\text{if}\;t\in[J_{n},J_{n+1}).$ (14)
which is generated by $\mathcal{A}$ with the transition function denoted as ${p(t,u,v)=e^{t\mathcal{A}}\mathbf{1}{v}(u)}{t\geq 0}$, where $\mathbf{1}{v}(u)$ represents the indicator function.
@@
See proofs in Appendix B.1. Theorem 5.1 is a variation of the well-known Feynman-Kac formula in stochastic analysis, which establishes a connection between a class of partial differential equations and stochastic differential equations (Freidlin, 1985). Based on the connection between the node feature $H(t)$ and the Markov chain $\mathbf{X}$, we can analyze the dynamics of diffusion-based GNNs by studying the properties of $\mathbf{X}$.
5.2 Limiting Distribution of $\mathbf{X}$
@@
$p(t,u,v)\rightarrow\mu(v),\quad t\rightarrow\infty.$ (16)
See proofs in Appendix B.2. The limit of probability measure $\mu$ in Eq.(16) is indeed an invariant measure of the generator $\mathcal{A}$. Due to its uniqueness, Theorem 5.2 shares the same measure $\mu$ as defined in Eq.(11). Theorem 5.2 further reflects the limiting states of the transition semigroup $\mathbf{P}=\{P_{t},t\geq 0\}$ corresponding to the generator $\mathcal{A}$. Combining with Theorem 5.1, we have
$h_{i}(u,t)$ $=\mathbb{E}^{u}[f_{i}(\mathbf{X}_{t})]=\sum_{v\in\mathcal{V}}f_{i}(v)p(t,u,v)$
$\rightarrow\sum_{v\in\mathcal{V}}X_{i}(v)\mu(v)\;\;\text{as}\;\;t\rightarrow\infty,$
showing that as the Markov chain transitions, each node feature tends to the average over the entire state space $\mathcal{V}$ of the input $X$. The uniqueness of such a limiting distribution stated in Theorem 5.2 indicates that each node feature will converge to the same value, providing the probabilistic interpretation to the oversmoothing issue in diffusion-based GNNs.
5.3 Killing Process
@@
$\tilde{\mathbf{X}}{t}:=\begin{cases}\mathbf{X}{t},\quad t<\tau\\ \partial,\quad t\geq\tau\end{cases}$ (17)
Here, $\tau$ is a stopping time, and $\partial$ signifies the dead state. We define the function $c(u)$ as:
$c(u):=-\lim\limits_{t\rightarrow 0^{+}}\frac{1}{t}\mathbb{P}(\tau\leq t|\mathbf{X}_{0}=u)$ (18)
The probability $\mathbb{P}(\tau\leq t|\mathbf{X}_{0}=u)$ represents the likelihood of being terminated, starting from state $u$ within the time interval $(0,t]$. Consequently, $c(u)$ can be comprehended as the negative of the killing rate. The following theorem asserts that the Markov process corresponding to the generator $\tilde{\mathcal{A}}=\mathcal{A}+c(u)$ indeed matches the killing process $\tilde{\mathbf{X}}$ defined in Eq. (17):
Theorem 5.3.
@@
(a) Effects of Training on Oversmoothing
@@
Figure 1: The left figure presents a comparative analysis of the Dirichlet energy for several diffusion-based GNNs, including GRAND-l, GRAND-nl, and Ours (Exp, Expn, Log, Sin, Cos), with (T) and without training, over increasing network depths; The right figure displays the effects of homophily level on oversmoothing. Error bands represent the variance in Dirichlet energy across different homophily levels, indicating each model’s sensitivity to homophily.
See proofs in Appendix B.3. In reference to Section 5.2, it has been established that the Markov process denoted as $\mathbf{X}$, characterized by the generator $\mathcal{A}$, will ultimately attain ergodicity across the entire state space. The probability measure $\mu(u)$, representing the likelihood of being in state $u\in\mathcal{V}$, remains unchanged, thereby causing different node features to converge towards a common value, which results in the oversmoothing phenomenon. However, the killing process $\tilde{\mathbf{X}}$ follows a distinct trajectory, as it transitions into a non-operational state, known as the “dead state,” after an unpredictable duration $\tau$. This behavior prevents it from achieving ergodicity throughout the state space and impedes the convergence of node features. It is worth noting that a similar concept is employed in the gating mechanisms frequently utilized in the field of deep learning. In fact, Rusch et al. (2023) has applied this gating mechanism to effectively address the oversmoothing problem.
5.4 Connection to Existing Work
@@
The probabilistic interpretation presented in this section can be seen as a broader extension of the insights derived from the discrete setting. Our research endeavors further delve into continuous-time Markov processes, with a particular emphasis on exploring the intricate relationships between stochastic processes, operator semigroups, and graph diffusion equations.
6 Experimental Verification

### v2

@@ 4.1. Operator Representation of Feature
(9) $h_{i}(u,t)=P_{t}f_{i}(u).$
The result in one dimension can naturally extend to multiple dimensions:
(10) $\begin{cases}h(u,t)=P_{t}f(u),\\ f(u)=X(u),\end{cases}\quad\text{for }u\in\mathcal{V},$
Discussion in section 4.2 shows the advantage of this form becomes, as we can derive specific form of the smoothing feature by ergodicity of operator.
4.2. Ergodicity and Oversmoothing
@@
(11) $\sum_{u\in\mathcal{V}}\mathcal{L}f(u)\pi(u)=0.$
We then define the ergodicity of generator $\mathcal{L}$ as follows:
Definition 2 (Ergodicity of $\mathcal{L}$).
The generator $\mathcal{L}$ is said to be ergodic if every $f$ such that $\mathcal{L}f=0$ is constant.
Next, we proceed to elaborate in the context of graph diffusion. It can be verified that the generator of graph diffusion $\mathcal{A}$ is ergodic, leading to the following theorem.
Theorem 3.
@@
A detailed proof is provided in Appendix A.2. From this theorem, it can be seen that the ergodicity of operator is dependent on the topology property of the graph. Notice that $10\leq a(u,v)<$ 1, from Perron-Frobenius theorem (Theorem 8.4.4 of (Horn and Johnson, 2012)), the invariant measure $\mu$ of $\mathcal{A}$ exists and is unique. Then the ergodicity of operator $\mathcal{A}$ ensures $P_{t}f(u)$ converges to the same constant for every $u\in\mathcal{V}$, as time $t\rightarrow\infty$. This finding is the cause of oversmoothing as different nodes share the same convergence. Under the same assumption in Theorem 3, above discussion can be summarized into the following theorem. Proof is provided in Appendix A.3.
Theorem 4.
@@
Remark 5.
From the proof of Theorem 4, we not only demonstrate that oversmoothing arises in diffusion-based GNNs but also that the convergence rate is associated with the spectral gap of $\mathcal{A}$. Furthermore, the results of Theorem 4 elucidate the specific mathematical form of the fixed point of oversmoothing, namely, the initial node features averaged over the invariant measure $\sum_{v\in\mathcal{V}}f(v)\mu(v)$. This constitutes an advancement over prior work, which only established the exponential contraction of certain smooth measures.
Since $P_{t}f$ converges to the eigenspace corresponding to the eigenvalue $\lambda_{0}=0$, once the generator is not ergodic, the projection of function $f\in\mathscr{L}^{2}(\mathcal{V},\mu)$ onto the eigenspace corresponding to the eigenvalue 00 is not constant. Therefore, $\lim\limits_{t\rightarrow\infty}P_{t}f$ is not constant. That means non-oversmoothing. Using the notation from Eq. (10) and combining with Theorem 4, we can derive the following convergence:
(13) $h(u,t)\rightarrow\sum_{v\in\mathcal{V}}X_{v}\mu(v).$
This provides a fundamental mathematical explanation for the oversmoothing in diffusion-based GNNs. The underlying reasons can be attributed to the ergodicity of the operator. Consequently, to relieve oversmoothing, it is natural to break ergodicity of operator, as shown in the section 4.3.
Further Discussion on Nonlinear Case. In addition to linear diffusion we further explore the more general case of over-smoothing in nonlinear diffusion models. We discuss the connection between weak ergodicity of operators and over-smoothing in nonlinear case.
The nonlinear diffusion-based GNNs is defined as follows
$\frac{\partial H(t)}{\partial t}:=[A(H(t))-I]H(t).$
After time discretization can be written as $H_{t+1}=A(H_{t})H_{t}=A(H_{t})\cdots A(H_{0})H_{0},\;t=0,1,2,\ldots$ where $H_{0}=X$. In this subsection we denote operator
(14) $P_{t}:=A(H_{t})\cdots A(H_{0}).$
For all $t=0,1,2,\ldots$, $A(H_{t})$ is a stochastic matrix. Thus the family of operators $\{P_{t}\},t=0,1,2,\ldots$ are the time inhomogeneous Markov operators, satisfies
(15) $h_{t}(u)=P_{t}f(u)=\sum_{v\in\mathcal{V}}P_{t}(u,v)f(v),\quad u\in\mathcal{V}$
where $f(u):=x(u)=h_{0}(u)$.
The research on the ergodicity of time inhomogeneous Markov operator is an important topic in probability (Douc et al., 2004; Saloff-Coste and Zúñiga, 2009). The following theorem shows the weak ergodicity of the operator $P_{t}$ in (14). Proof is provided in Appendix A.4.
Theorem 6.
If The graph $\mathcal{G}$ is connected, non-bipartite and have self-loop on each node, assume the attention function $\Phi$ is continuous. Then Markov operator $P_{t}$ is weak ergodic, that is
(16) $\lim\limits_{t\to\infty}\|P_{t}(u,\cdot)-P_{t}(v,\cdot)\|_{\mathrm{TV}}=0,\quad\forall u,v\in\mathcal{V},$
where $\|\cdot\|_{\mathrm{TV}}$ is total variation distance.
From Theorem 6 we have for all $u,v\in\mathcal{V}$,
$\lim\limits_{t\to\infty}h_{t}(u)$ $=\lim\limits_{t\to\infty}P_{t}f(u)=\lim\limits_{t\to\infty}\sum_{w\in\mathcal{V}}P_{t}(u,w)f(w)$
$=\lim\limits_{t\to\infty}\sum_{w\in\mathcal{V}}P_{t}(v,w)f(w)=\lim\limits_{t\to\infty}h_{t}(v).$
This shows that nonlinear diffusion-based GNNs still suffer from oversmoothing.
4.3. Ergodicity Breaking Term
We aim to relieve oversmoothing by breaking the ergodicity of the operator. This necessitates a modification of the operator $\mathcal{A}$ as presented in Eq. (7). The central concept underpinning our approach involves introducing an additional term to the operator $\mathcal{A}$, yielding a new operator:
(17) $\tilde{\mathcal{A}}:=\mathcal{A}+\mathcal{C},$
@@ 4.3. Ergodicity Breaking Term
Proof for Theorem 7 is furnished in Appendix A.5. By substituting the operator $\mathcal{A}$ with the operator $\tilde{\mathcal{A}}$, we obtain a variation of the diffusion process (4): $\frac{\partial h(u,t)}{\partial t}:=\mathcal{A}h(u,t)+\mathcal{C}h(u,t),$ serving as a framework for designing GNNs to relieve oversmoothing.
5. Probability View for Oversmoothing
@@
5.1. Markov Chain on Graph with Generator $\mathcal{A}$
Given a Markov process denoted as $\xi_{t}$ characterized by its state space $\mathscr{E}$ and a transition kernel denoted as $p(t,x,\mathrm{d}y)$, we can naturally establish a Markov semigroup as: $P_{t}f(x):=\int_{E}f(y)p(t,x,\mathrm{d}y)=\mathbb{E}^{x}[f(\xi_{t})],$ where $\mathbb{E}^{x}[\cdot]:=\mathbb{E}[\cdot|\xi_{0}=x]$ represents the expectation operator under the probability measure $p(\cdot,x,\mathrm{d}y)$. The generator of the Markov process $\xi_{t}$ can be obtained via Equation (6). Conversely, in accordance with the Kolmogorov theorem (Theorem 2.11 of (Ge and Blumenthal, 2011)), there exists a Markov process given its transition function and initial distribution. In this subsection, commencing from the generator $\mathcal{A}$ of graph diffusion, we formulate the corresponding process denoted as $\mathbf{X}$, which serves as a representation for the node features $H(t)$.
Let $(\Omega,\mathscr{F},\mathbb{P})$ be a probability space and $\mathbf{Y}=\{\mathbf{Y}_{n}\}_{n\in\mathbb{N}}$ be a Markov chain with state space $(\mathcal{V},\mathscr{S})$ such that $\mathbb{P}(\mathbf{Y}_{n+1}=v|\mathbf{Y}_{n}=u)=a(u,v).$ The jumping times are defined as:
@@ 5.1. Markov Chain on Graph with Generator $\mathcal{A}$
where $\zeta_{k},\;k=1,2,\ldots,n$ is a sequence of independent exponentially distributed random variables of parameter $1$.
Next, we introduce the right-continuous process $\mathbf{X}:[0,\infty)\times\Omega\rightarrow X$ defined as:
(18) $\mathbf{X}_{t}:=\mathbf{Y}_{n},\;\text{if}\;t\in[J_{n},J_{n+1}).$
which is generated by $\mathcal{A}$ with the transition function denoted as ${p(t,u,v)=e^{t\mathcal{A}}\mathbf{1}_{v}(u)}{t\geq 0}$, where $\mathbf{1}_{v}(u)$ represents the indicator function.
@@
Theorem 1 is a variation of the well-known Feynman-Kac formula in stochastic analysis, which establishes a connection between a class of partial differential equations and stochastic differential equations (Freidlin, 1985). Based on the connection between the node feature $H(t)$ and the Markov chain $\mathbf{X}$, we can analyze the dynamics of diffusion-based GNNs by studying the properties of $\mathbf{X}$.
5.2. Limiting Distribution of Node Features
@@
(20) $p(t,u,v)\rightarrow\mu(v),\quad t\rightarrow\infty.$
The limit of probability measure $\mu$ in Eq.(20) is indeed an invariant measure of the generator $\mathcal{A}$. Due to its uniqueness, Theorem 2 shares the same measure $\mu$ as defined in Eq.(12). Theorem 2 further reflects the limiting states of the transition semigroup $\mathbf{P}=\{P_{t},t\geq 0\}$ corresponding to the generator $\mathcal{A}$. Combining with Theorem 1, we have
$h_{i}(u,t)$ $=\mathbb{E}^{u}[f_{i}(\mathbf{X}_{t})]=\sum_{v\in\mathcal{V}}f_{i}(v)p(t,u,v)\rightarrow\sum_{v\in\mathcal{V}}X_{i}(v)\mu(v)\;\;$
as $t\rightarrow\infty,$ showing that as the Markov chain transitions, each node feature tends to the average over the entire state space $\mathcal{V}$ of the input $X$. The uniqueness of such a limiting distribution stated in Theorem 2 indicates that each node feature will converge to the same value, providing the probabilistic interpretation to the oversmoothing issue in diffusion-based GNNs.
5.3. Killing Process
@@
(21) $\tilde{\mathbf{X}}_{t}:=\begin{cases}\mathbf{X}_{t},\quad t<\tau\\ \partial,\quad t\geq\tau\end{cases}$
Here, $\tau$ is a stopping time, and $\partial$ signifies the dead state. We define the function $c(u)$ as: $c(u):=-\lim\limits_{t\rightarrow 0^{+}}\frac{1}{t}\mathbb{P}(\tau\leq t|\mathbf{X}_{0}=u)$. The probability $\mathbb{P}(\tau\leq t|\mathbf{X}_{0}=u)$ represents the likelihood of being terminated, starting from state $u$ within the time interval $(0,t]$. Consequently, $c(u)$ can be comprehended as the negative of the killing rate. The following theorem asserts that the Markov process corresponding to the generator $\tilde{\mathcal{A}}=\mathcal{A}+c(u)$ indeed matches the killing process $\tilde{\mathbf{X}}$ defined in Eq. (21):
Theorem 3.
@@
(a) Oversmoothing on Wisconsin Dataset
@@
(c) Performance as Layer Increases
Figure 2. Demonstration on Oversmoothing and Performance Metrics. The figure comprises three panels: 2(a) showcases the increase in oversmoothing of various GNN models with increasing depth on the Wisconsin dataset, with models like GCN and GAT experiencing significant effects; 2(b) illustrates the impact of synthetic Cora dataset with varying homophily levels on oversmoothing; and 2(c) displays the corresponding decrease in test accuracy on Cora dataset as network depth increases, highlighting performance declines in models such as GCN and GAT at greater depths due to the oversmoothing effects noted in the other panels.
In reference to Section 5.2, it has been established that the Markov process denoted as $\mathbf{X}$, characterized by the generator $\mathcal{A}$, will ultimately attain ergodicity across the entire state space. The probability measure $\mu(u)$, representing the likelihood of being in state $u\in\mathcal{V}$, remains unchanged, thereby causing different node features to converge towards a common value, which results in the oversmoothing phenomenon. However, the killing process $\tilde{\mathbf{X}}$ follows a distinct trajectory, as it transitions into a non-operational state, known as the “dead state,” after an unpredictable duration $\tau$. This behavior prevents it from achieving ergodicity throughout the state space and impedes the convergence of node features. It is worth noting that a similar concept is employed in the gating mechanisms frequently utilized in the field of deep learning. In fact, Rusch et al. (2023) has applied this gating mechanism to effectively address the oversmoothing problem.
6. Connections and Extensions to Existing Works
In this section, we explore the relationships between this work and the existing literature, highlighting both the continuities and the novelties introduced.
6.1. Confirmation of Oversmoothing
@@
The probabilistic interpretation presented in section 5 can be seen as a broader extension of the insights derived from the discrete setting. Our research endeavors further delve into continuous-time Markov processes, with a particular emphasis on exploring the intricate relationships between stochastic processes, operator semigroups, and graph diffusion equations.
6.2. Relief of Oversmoothing
Graph diffusion with ergodicity breaking term $\mathcal{C}$ proposed in Sec. 4.3 is an explanatory framework for many existing diffusion-based GNNs that have been proven effective in addressing the oversmoothing problem. In previous attempts to mitigate the oversmoothing issue in diffusion-based GNNs, researchers have proposed several approaches. Chamberlain et al. (2021b) adds a source term in the graph diffusion equation, while Wang et al. (2023) alleviates the issue of oversmoothing by introducing an Allen-Cahn term. Based on the reaction-diffusion equations, Choi et al. (2023) introduces the reaction term $r(H(t))$, encompassing the forms of the previous two approaches. In comparison to the aforementioned methods (Thorpe et al., 2022; Choi et al., 2023; Wang et al., 2023), our framework offers a more rigorous theoretical foundation and provides the essential condition to deal with oversmoothing, that is, breaking ergodicity. Then the three methods mentioned earlier, which address the oversmoothing issue by incorporating additional terms, can be viewed as specific cases of our approach.
6.3. Technical Improvements
The operator semigroup method serves as a robust tool for analyzing both linear and nonlinear dynamic systems. In this work, we do not endeavor to prove the exponential contraction of a smooth measure (Nguyen et al., 2023; Wu et al., 2023). Instead, utilizing the semigroup method, we conduct a thorough analysis of the dynamics of the graph diffusion equation across varying initial conditions. By employing definitions of ergodicity and invariant measures associated with operators, we establish a specific mathematical expression for the smoothing feature, advancing beyond previous studies that merely confirmed the existence of oversmoothing. Additionally, the operator semigroup framework facilitates our investigation into the quantitative bounds of convergence rates through functional inequalities (Bakry et al., 2014) in future work. Furthermore, our approach fosters a deep connection between probability theory and functional analysis via operator semigroups and generators, offering a unified perspective that encompasses many earlier works framed within the context of Markov processes (Thorpe et al., 2022). Graph signal processing (GSP) emerged as one of the pioneering frameworks for elucidating the phenomenon of oversmoothing (Li et al., 2018). Within this framework, the graph convolution operator functions analogously to a low-pass filter, resulting in smoothing of node features. GSP is grounded in spectral graph theory, which emphasizes operators associated with the graph’s structure such as the Laplacian. This foundation aligns with operator semigroups, which focus on operators that act on node features. While, semigroup methods show more power in continuous models.
7. Experimental Verification